# Planar site percolation via tree embeddings

### v5

@@ Abstract.
Let $G$ be an infinite, connected planar graph that is properly embedded in ${\mathbb{R}}^{2}$ and has minimal vertex degree at least $7$. We prove that for i.i.d. $\mathrm{Bernoulli}(p)$ site percolation on $G$, for every
@@
there are almost surely infinitely many infinite open $(1$-) clusters. In particular, we show that $p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2}$, so the non-uniqueness interval is nonempty. This verifies Conjecture 7 of Benjamini and Schramm [5] for this class of properly embedded planar graphs.
The main new ingredient is an explicit construction of trees embedded in $G$ with controlled local geometry (degrees $3$ and $4$ away from the root). These trees yield non-uniqueness near $p=\tfrac{1}{2}$ and, via an embedded forest of pairwise disjoint trees, imply exponential decay of two-point connection probabilities in the matching graph $G_{*}$. To treat the complementary regime $p>\tfrac{1}{2}$, we introduce a “binary-tree” version of uniform percolation that yields stability of infinite clusters under edge additions without any bounded-degree assumption.
Beyond planarity, we also obtain a general two-point connectivity inequality on arbitrary infinite, connected locally finite graphs, which implies a non-transitive criterion for the existence of infinitely many infinite clusters in the regime $p_{c}<p<p_{\mathrm{conn}}$.
@@ 1.1. Overview and Main Results
Introduced by Broadbent and Hammersley [9], percolation is a fundamental model of random substructure on graphs, exhibiting a sharp phase transition between the absence and presence of an infinite connected component; see Grimmett [16] and the survey of Duminil-Copin [12] for background and recent developments.
Let $G=(V,E)$ be a graph. We write $e=\langle u,v\rangle$ for an edge with endpoints $u,v\in V$. The (vertex-)degree of a vertex $v\in V$ is the number of edges incident to $v$. We say that $G$ is locally finite if every vertex has finite degree.
Assume throughout that $G=(V,E)$ is infinite, locally finite and connected. A (site) percolation configuration is $\omega\in\{0,1\}^{V}$, where $\omega(v)=1$ means that $v$ is open and $\omega(v)=0$ means that $v$ is closed. A cluster is a maximal connected set of vertices with the same state. We call a cluster a 1-cluster (resp. 0-cluster) if all of its vertices are open (resp. closed). A cluster may be finite or infinite, according to its number of vertices.
In this paper, we consider planar graphs admitting proper embeddings in the plane.
Definition 1.1.
A planar graph $G$ is a graph that can be drawn in the plane ${\mathbb{R}}^{2}$, with vertices represented by points and edges represented by curves, such that edges intersect only at their endpoints.
We say a graph is properly embedded in ${\mathbb{R}}^{2}$ if it can be drawn in the plane such that any compact subset of ${\mathbb{R}}^{2}$ intersects only finitely many vertices and edges.
See also Section 2 of [7]. In particular, a properly embedded planar graph is locally finite. Since ${\mathbb{R}}^{2}$ is homeomorphic to any simply-connected open subset of ${\mathbb{R}}^{2}$, a graph can be properly embedded into ${\mathbb{R}}^{2}$ if and only if it can be properly embedded into any simply-connected open subset of ${\mathbb{R}}^{2}$ (for instance, the unit disk; with a suitable metric this leads to embeddings in the hyperbolic plane ${\mathbb{H}}^{2}$). We refer to [11, 7, 30, 17, 18, 19] for related work on graphs in ${\mathbb{H}}^{2}$ and statistical mechanical models. We also note that a graph generally cannot be circle packed in both ${\mathbb{R}}^{2}$ and ${\mathbb{H}}^{2}$; see [24, 25].
Of particular interest is i.i.d. Bernoulli site percolation. In this model, $\{\omega(v)\}_{v\in V}$ are independent Bernoulli random variables with ${\mathbb{P}}_{p}(\omega(v)=1)=p\in[0,1]$. We write ${\mathbb{P}}_{p}$ for the associated product measure.
For i.i.d. Bernoulli site percolation, define
(1.1) $p_{c}^{site}(G)$ $:=\inf\Bigl\{p\in[0,1]:{\mathbb{P}}_{p}\bigl(\text{$G$ has an infinite 1-cluster}\bigr)=1\Bigr\},$
(1.2) $p_{u}^{site}(G)$ $:=\inf\Bigl\{p\in[0,1]:{\mathbb{P}}_{p}\bigl(\text{$G$ has a unique infinite 1-cluster}\bigr)=1\Bigr\}.$
Clearly $p_{c}^{site}(G)\leq p_{u}^{site}(G)$. When the inequality is strict, there is a non-uniqueness phase in which (with positive probability) there exist at least two infinite 1-clusters.
A number of problems on uniqueness and non-uniqueness of infinite clusters were formulated by Benjamini and Schramm in their influential paper [5], including the following.
@@
(Conjecture 7 in [5]). Consider site percolation on an infinite, connected, planar graph $G$ with minimal degree at least $7$. Then for any $p\in(p_{c}^{site}(G),\,1-p_{c}^{site}(G))$ we have ${\mathbb{P}}_{p}$-a.s. there are infinitely many infinite 1-clusters in the i.i.d. Bernoulli($p$) site percolation on $G$. Moreover, $p_{c}^{site}(G)<\tfrac{1}{2}$, so the above interval is non-empty.
Context. The number of infinite clusters in supercritical percolation is classical and well understood on transitive (and, more generally, quasi-transitive) graphs, where symmetry-based tools such as 0–1 laws and the mass-transport principle apply; see e.g. [3, 4]. Aizenman–Kesten–Newman [2] proved that for i.i.d. Bernoulli percolation on a transitive graph the number of infinite clusters can only be $0,1$ or $\infty$. Burton–Keane [10] excluded the case $\infty$ for ${\mathbb{Z}}^{d}$, and Gandolfi–Keane–Newman [15] extended this exclusion to quasi-transitive amenable graphs.
A closely related conjecture in [5] concerns non-amenable quasi-transitive graphs.
Conjecture 1.3.
(Conjecture 6 in [5]) Assume that $G$ is quasi-transitive and non-amenable. Then $p_{c}(G)<p_{u}(G)$.
There is substantial progress on Conjecture 1.3 under additional assumptions; see, for example, [20, 37, 42, 14, 36, 34, 26, 27]. Despite these advances, Conjecture 1.3 remains open.
Planarity introduces additional structure. In the quasi-transitive planar setting, non-uniqueness can be proved using duality/matching arguments: Lalley studied percolation on Fuchsian groups [29], Benjamini–Schramm studied percolation in the hyperbolic plane [7], and Lyons–Peres [35] extended the identity $p_{c}^{bond}(G)+p_{u}^{bond}(G^{+})=1$ to the quasi-transitive case. Via the correspondence between bond configurations on dual pairs and site configurations on matching pairs, one obtains $p_{c}^{site}(G)+p_{u}^{site}(G_{*})=1$ in many quasi-transitive planar situations; see [30, 18, 19]. This confirms Conjecture 1.2 in the quasi-transitive case.
For general planar graphs without quasi-transitivity, these symmetry-based tools are unavailable. Haslegrave and Panagiotis [23] proved the inequality $p_{c}^{site}(G)<\tfrac{1}{2}$ without any quasi-transitivity assumption, but the main conclusion of Conjecture 1.2—the existence of infinitely many infinite clusters throughout $(p_{c}^{site},\,1-p_{c}^{site})$—remained open.
Main theorem. In this paper we verify Conjecture 1.2 for properly embedded planar graphs.
@@
Let $G$ be an infinite, connected graph properly embedded in ${\mathbb{R}}^{2}$ with minimal vertex degree at least $7$. Then
  • $p_{c}^{site}(G)<\tfrac{1}{2}$;
@@
  • for any $\in(p_{c}^{site}(G),\,1-p_{c}^{site}(G)), p{\mathbb{P}}_{p}$-a.s. there are infinitely many infinite 1-clusters in i.i.d. Bernoulli($p$) site percolation on $G$.
New ingredients and proof strategy. The proof rests on an explicit construction of trees embedded in $G$ with degrees $3$ and $4$ away from the root. This construction is more geometric than the spanning-forest existence result of [6], and it is tailored to provide separation in the plane.
A second key role is played by the matching graph, which is the site-percolation analogue of the dual graph in bond percolation.
Definition 1.5.
  1. (1)
    $\langle u,v\rangle\in E$; or
  2. (2)
    $\langle u,v\rangle\notin E$, but $u$ and $v$ share a finite face in $G$.
The graphs considered in this paper may have more than one end; hence we allow infinite faces. In Definition 1.5, the additional edges $E_{*}\setminus E$ only join vertices sharing finite faces, so that $G_{*}$ remains locally finite. If every face of a proper embedding of $G$ is finite, then Definition 1.5 coincides with the usual definition of matching graphs.
We also recall the notion of ends.
Definition 1.6.
At a high level, the proof of Theorem 1.4 proceeds by combining (i) tree embeddings that give non-uniqueness around $p=\tfrac{1}{2}$, (ii) an embedded forest that yields exponential decay of two-point connectivities in $G_{*}$ when $p<1-p_{c}^{site}(T)$, and (iii) a “binary-tree” uniform percolation mechanism that provides stability of infinite clusters under edge additions for $p>\tfrac{1}{2}$. These ingredients allow us to cover the entire interval $(p_{c}^{site}(G),\,1-p_{c}^{site}(G))$.
A general criterion beyond planarity. Although our main theorems concern planar graphs, part of the underlying mechanism is purely graph-theoretic and does not rely on planarity, transitivity, or any bounded-degree assumption. We record here a general connectivity criterion, which may be of independent interest.
Definition 1.7 (Connectivity-decay threshold).
Let $G=(V,E)$ be an infinite, connected, locally finite graph. Define
$p_{\mathrm{conn}}(G):=\sup\Bigl\{p\in(0,1):\ \lim_{n\to\infty}\ \sup_{\begin{subarray}{c}x,y\in V\\ d_{G}(x,y)\geq n\end{subarray}}\mathbb{P}_{p}(x\leftrightarrow y)=0\Bigr\},$
where $\mathbb{P}_{p}$ denotes i.i.d. $\mathrm{Bernoulli}(p)$ site percolation on $G$.
Let $\mathcal{A}_{1}$ be the event that there exists a unique infinite open $(1$-) cluster.
Proposition 1.8 (Two-point lower bound under uniqueness).
Let $G=(V,E)$ be an infinite, connected, locally finite graph. For every $u,v\in V$,
@@
Corollary 1.9 (Infinitely many infinite clusters below $p_{\mathrm{conn}}$).
Let $G=(V,E)$ be an infinite, connected, locally finite graph. If $p_{c}^{\mathrm{site}}(G)<p<p_{\mathrm{conn}}(G)$, then $\mathbb{P}_{p}$-almost surely there exist infinitely many infinite open clusters.
In the planar setting, our tree-embedding and embedded-forest constructions provide quantitative decay of two-point connectivities in suitable matching graphs, which can be combined with Corollary 1.9 to force non-uniqueness in the complementary supercritical regime.
Further applications and companion results. Variants of these ideas were later used in [32] to construct a counterexample to Conjecture 7 of [5] for planar graphs with uncountably many ends. Finally, the methods developed here have further applications: in [31] they are used to prove a vertex-cut characterization of $p_{c}^{\mathrm{site}}$ (conjectured by Kahn in [28]) and to refute an edge-cut characterization proposed by Lyons and Peres [35] and Tang [43].
Organization of the paper.
  1. (1)
    In Section 2 we explicitly construct an embedded tree $T\subseteq G$ and show $p_{c}^{site}(G)\leq p_{c}^{site}(T)<\tfrac{1}{2}$; moreover, at $p=\tfrac{1}{2}$ there are a.s. infinitely many infinite 1-clusters.
  2. (2)
    In Section 4 we construct an embedded forest of disjoint trees that separates pairs of vertices and yields exponential decay of two-point connection probabilities in $G_{*}$ for $p<1-p_{c}^{site}(T)$.
  3. (3)
    In Section 3 we combine the exponential decay with a general connectivity criterion (Section 3) to deduce infinitely many infinite clusters in the subcritical-to-intermediate regime.
  4. (4)
    Sections 56 treat the near-$1$ regime using planar duality and the end structure of 0-*-clusters, together with bounds on non-self-touching polygons (verified first for triangulations).
  5. (5)
    Section 7 introduces a binary-tree version of uniform percolation and proves stability of infinite clusters under adding edges, completing the proof for general properly embedded planar graphs.
Two methodological by-products may be of independent interest. First, Step (3) relies on a characterization of $p_{c}^{site}$ for all infinite, connected, locally finite graphs without any transitivity assumption, extending arguments of [22, 13]. Second, the binary-tree uniform percolation in Step (5) avoids bounded-degree assumptions in stability arguments inspired by [41].
1.2. Background and Notation
@@ 2. Embedded Trees
Step 1: $p_{T}<\tfrac{1}{2}$. By [33, Thm. 6.2] (see also [35, Thm. 5.15]), for any infinite locally finite tree,
@@
Figure $2.$ Infinite open clusters in the tree rooted at $v_{r01}$ is separated from the infinite open clusters in the tree rooted at $v_{r11}$ by the infinite closed cluster occupying $v_{r}$, $v_{r0},v_{r1}$, $v_{r00},v_{r10}$
@@
If we replace the assumption ‘the minimal vertex degree is at least 7” in Lemmas 2.10 and 2.11 by the minimal vertex degree is at least 5; and the minimal face degree is at least 4”, the same conclusion can be proved similarly.

### v6

# Tree embeddings and nonuniqueness in site percolation

@@ Abstract.
We prove a nonuniqueness theorem for Bernoulli site percolation on properly embedded planar graphs, and we obtain a general connectivity principle beyond planarity. Let $G$ be an infinite connected graph properly embedded in ${\mathbb{R}}^{2}$ with minimum degree at least $7$. Then
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2},$
and for every
@@
Bernoulli(p)(p) site percolation on $G$ has almost surely infinitely many infinite open clusters. In particular, this verifies a conjecture of Benjamini and Schramm for properly embedded planar graphs.
The core new ingredient is an explicit embedded-tree separation mechanism for planar nonuniqueness. We construct embedded trees and an embedded forest whose separation properties yield exponential decay of two-point connection probabilities in the matching graph. To treat the high-density regime, we introduce a binary-tree version of uniform percolation and prove stability of infinite clusters under edge additions, without any bounded-degree assumption.
Beyond the planar theorem, we prove a general lower bound on two-point connectivity under uniqueness for arbitrary infinite locally finite graphs. As a consequence, if
$p_{c}^{\mathrm{site}}(G)<p<p_{\mathrm{conn}}(G),$
then Bernoulli site percolation on $G$ has almost surely infinitely many infinite open clusters.
@@ 1.1. Overview and Main Results
A central problem in percolation theory is to understand the number of infinite clusters in the supercritical phase. On transitive graphs this question is by now classical: for Bernoulli percolation, Aizenman, Kesten and Newman [2] proved that the number of infinite clusters is almost surely 0, 11, or $\infty$, while Burton and Keane [7] and Gandolfi, Keane and Newman [10] established uniqueness throughout the supercritical phase for large classes of amenable graphs. In contrast, on non-amenable graphs one expects a genuine nonuniqueness phase, and a central conjecture of Benjamini and Schramm [3] predicts that $p_{c}<p_{u}$ for every non-amenable quasi-transitive graph.
Planarity offers a different route to nonuniqueness, through duality and matching-graph ideas. In quasi-transitive planar graphs these methods lead to a rather complete picture; see, for example, [5, 22, 18, 12, 13]. The situation is much less understood without quasi-transitivity, where symmetry-based tools such as mass transport are unavailable. The purpose of this paper is to show that, in this nonsymmetric planar setting, one can nevertheless prove nonuniqueness throughout an entire interval by a purely geometric argument.
We consider i.i.d. Bernoulli site percolation on infinite connected planar graphs admitting proper embeddings in ${\mathbb{R}}^{2}$, that is, embeddings for which every compact subset of ${\mathbb{R}}^{2}$ intersects only finitely many vertices and edges. For such a graph $G$, write $p_{c}^{\mathrm{site}}(G)$ and $p_{u}^{\mathrm{site}}(G)$ for the critical and uniqueness thresholds:
(1.1) $p_{c}^{\mathrm{site}}(G)$ $:=\inf\Bigl\{p\in[0,1]:{\mathbb{P}}_{p}\bigl(\text{$G$ has an infinite open cluster}\bigr)=1\Bigr\},$
(1.2) $p_{u}^{\mathrm{site}}(G)$ $:=\inf\Bigl\{p\in[0,1]:{\mathbb{P}}_{p}\bigl(\text{$G$ has a unique infinite open cluster}\bigr)>0\Bigr\}.$
Clearly $p_{c}^{\mathrm{site}}(G)\leq p_{u}^{\mathrm{site}}(G)$. When the inequality is strict, there is a nonuniqueness phase in which infinitely many infinite clusters are expected to appear.
Benjamini and Schramm [3, Conjecture 7] proposed the following planar degree-$7$ picture.
@@
Let $G$ be an infinite connected planar graph with minimum degree at least $7$. Then
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2},$
and for every
$p\in\bigl(p_{c}^{\mathrm{site}}(G),\,1-p_{c}^{\mathrm{site}}(G)\bigr),$
Bernoulli site percolation on $G$ has almost surely infinitely many infinite open clusters.
In quasi-transitive planar graphs, nonuniqueness can be recovered from matching-graph duality and symmetry arguments. Without quasi-transitivity, however, these tools are no longer available. Haslegrave and Panagiotou [16] proved that
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2}$
under the minimum-degree assumption and the proper embedding assumption, but the corresponding nonuniqueness statement remained open in this level of generality.
Our main result establishes the full conjectural picture for properly embedded planar graphs.
@@
Let $G$ be an infinite connected graph properly embedded in ${\mathbb{R}}^{2}$, and assume that every vertex of $G$ has degree at least $7$. Then
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2}.$
Moreover, for every
$p\in\bigl(p_{c}^{\mathrm{site}}(G),\,1-p_{c}^{\mathrm{site}}(G)\bigr),$
@@
i.i.d. Bernoulli(p)(p) site percolation on $G$ has almost surely infinitely many infinite open clusters.
The point of Theorem 1.2 is conceptual as well as quantitative. It shows that planar nonuniqueness can be proved in the absence of transitivity by an explicit geometric mechanism. In particular, the proof does not rely on quasi-transitivity, mass transport, or bounded-degree assumptions.
1.2. The Geometric Setting
The geometric content of the problem enters through the proper embedding. In the quasi-transitive planar setting, duality and matching-graph arguments can often be used at a global level and then converted, via symmetry, into statements about the number of infinite clusters. In the present setting there is no such symmetry, so the embedding itself must supply the mechanism that replaces it.
Proper embeddings play two roles in our argument. First, they allow us to work with finite faces in a locally finite planar environment and to pass naturally to the matching graph. Informally, the matching graph $G_{*}$ is obtained from $G$ by adding edges between vertices that lie on a common finite face. It is the site-percolation analogue of the planar dual, but here it must be used quantitatively rather than through symmetry identities alone.
Second, proper embeddings make it possible to construct explicit separating objects in the plane. The embedded trees and forests built in this paper are not merely auxiliary subgraphs: they provide the geometric barriers that force long connections to cross many disjoint regions. This separation mechanism is the source of the quantitative decay estimates that later drive the nonuniqueness argument.
A further point is that the graphs considered here need not be quasi-transitive and may have nontrivial end structure. In particular, one cannot reduce the analysis to a homogeneous large-scale geometry. The proof therefore has to extract enough rigidity directly from the embedding and from the degree assumption.
1.3. Proof Strategy
We now describe the architecture of the proof of Theorem 1.2. The argument proceeds in three steps, each addressing a different obstruction to nonuniqueness.
The first step is to build an explicit embedded tree $T\subseteq G$ with controlled local geometry. This tree already captures enough expansion to imply
$p_{c}^{\mathrm{site}}(G)\leq p_{c}^{\mathrm{site}}(T)<\tfrac{1}{2}.$
Moreover, the tree structure yields nonuniqueness near $p=\tfrac{1}{2}$. The point is not only to compare critical parameters, but to produce a rigid geometric object whose branches can later be organized into separating families.
Our embedded-tree construction should be compared with the spanning-forest theorem of Benjamini and Schramm [4], which shows that graphs with positive Cheeger constant contain trees with positive Cheeger constant. The present construction is of a different nature: it is explicit, adapted to the planar embedding, and designed to produce left/right separation in the plane. This additional geometric control is what later yields a disjoint embedded forest and exponential decay of two-point connection probabilities in the matching graph.
The second step is to upgrade this construction to an embedded forest of pairwise disjoint trees. The role of this forest is to separate distant vertices in a quantitative way. Roughly speaking, any long connection in the matching graph must cross many disjoint barriers created by the forest, and this leads to exponential decay of two-point connection probabilities in $G_{*}$. This gives control of the part of the supercritical regime in which long-range connectivity in the matching graph is still sufficiently weak.
The third step treats the complementary high-density regime. At this point one needs to understand how infinite clusters behave under the addition of edges, and this becomes delicate without a bounded-degree assumption. To overcome this difficulty we introduce a binary-tree version of uniform percolation and prove a stability statement for infinite clusters under edge additions. This allows us to transfer nonuniqueness information from the matching graph back to the original graph.
A key probabilistic input enters through a general lower bound on two-point connectivity under uniqueness. Let $\mathcal{A}_{1}$ denote the event that there exists a unique infinite open cluster. We prove the following statement, valid for arbitrary infinite connected locally finite graphs.
Proposition 1.3.
Let $G=(V,E)$ be an infinite, connected, locally finite graph. Then for every $u,v\in V$,
@@ 1.3. Proof Strategy
This yields a general criterion for nonuniqueness. Define
(1.3) $p_{\mathrm{conn}}(G):=\sup\Bigl\{p\in(0,1):\ \lim_{n\to\infty}\sup_{\begin{subarray}{c}x,y\in V\\ d_{G}(x,y)\geq n\end{subarray}}{\mathbb{P}}_{p}(x\leftrightarrow y)=0\Bigr\}.$
Corollary 1.4.
Let $G=(V,E)$ be an infinite, connected, locally finite graph. If
$p_{c}^{\mathrm{site}}(G)<p<p_{\mathrm{conn}}(G),$
then Bernoulli site percolation on $G$ has almost surely infinitely many infinite open clusters.
The criterion in Corollary 1.4 is completely independent of planarity and transitivity and may be viewed as a general probabilistic principle: uniqueness of the infinite cluster forces a nontrivial lower bound on long-range two-point connectivity. In the planar setting, the embedded-forest construction supplies the required decay estimate in the matching graph, and Corollary 1.4 converts this quantitative input into nonuniqueness. Theorem 1.2 then follows by combining this with the stability mechanism described above.
1.4. Relation to Previous and Concurrent Work
On quasi-transitive planar graphs, nonuniqueness phenomena can often be analyzed through duality and matching-graph identities; see, for example, [5, 22, 18, 12, 13]. The present paper addresses a different regime, where quasi-transitivity is absent and the main issue is to replace symmetry-based tools by explicit geometric constructions.
For properly embedded planar graphs with minimum degree at least $7$, Haslegrave and Panagiotou [16] proved the inequality
(1.4) $p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2}.$
Our contribution is a new proof of (1.4), and to establish the corresponding nonuniqueness statement throughout
$\bigl(p_{c}^{\mathrm{site}}(G),\,1-p_{c}^{\mathrm{site}}(G)\bigr)$
for this class of graphs.
After earlier versions of this work, Glazman, Harel and Zelesko [11] proved a general $0/\infty$ theorem for planar percolation processes, covering Bernoulli site percolation at $p\leq\tfrac{1}{2}$. Subsequent work [20] shows that, in full generality, Conjecture 7 is false, although under a natural countability assumption on end-equivalence classes one recovers the same nonuniqueness interval. Against this background, the main novelty of the present paper is to identify an explicit geometric route to planar nonuniqueness without transitivity, based on embedded trees, planar separation, and stability under edge additions.
1.5. Organization of the Paper
Section 2 collects background material and notation. In Section 3 we construct the embedded tree and prove
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2},$
together with nonuniqueness in the interval
$\bigl(p_{c}^{\mathrm{site}}(T),\,1-p_{c}^{\mathrm{site}}(T)\bigr).$
Section 4 develops the general connectivity criterion for arbitrary locally finite graphs, including the lower bound on two-point connectivity under uniqueness and its nonuniqueness consequence. In Section 5 we build the embedded-forest separation scheme and derive exponential decay of two-point connection probabilities in the matching graph; combined with the results of Section 4, this yields nonuniqueness in the intermediate regime. Section 6 reduces the near-critical high-density regime to a polygon-counting hypothesis by analyzing end structure and infinite 0-*-clusters. Section 7 verifies this hypothesis for triangulations. Finally, Section 8 introduces the binary-tree version of uniform percolation and a triangulation/stability argument, completing the proof of Theorem 1.2 for general properly embedded planar graphs.
2. Background and Notation
@@
Definition 2.1.
  1. (1)
    $\langle u,v\rangle\in E$; or
  2. (2)
    $\langle u,v\rangle\notin E$, but $u$ and $v$ share a finite face in $G$.
Definition 2.2.
@@ 3. Embedded Trees
Step 1: $p_{T}<\tfrac{1}{2}$. By [21, Thm. 6.2] (see also [22, Thm. 5.15]), for any infinite locally finite tree,
@@
Figure 3.2. Infinite open clusters in the tree rooted at $v_{r01}$ is separated from the infinite open clusters in the tree rooted at $v_{r11}$ by the infinite closed cluster occupying $v_{r}$, $v_{r0},v_{r1}$, $v_{r00},v_{r10}$
@@
If we replace the assumption ‘the minimal vertex degree is at least 7” in Lemmas 3.10 and 3.11 by the minimal vertex degree is at least 5; and the minimal face degree is at least 4”, the same conclusion can be proved similarly.